Suppose that there are two very large reservoirs of water, one at a temperature of 95.0 °C, and one at a temperature of 11.0 ℃. These reservoirs are brought into thermal contact long enough for 36330 J of heat to flow from the hot water to the cold water. Assume that the reservoirs are large enough so that the temperatures do not change significantly What is the total change in entropy resulting from this heat exchange between the hot water and the cold water? Number J/K Calculate the amount of energy made unavailable for work by this increase in entropy. Number How much work could a Canot engine do if it took in the given amount of heat (36330 J) from the hot water reservoir, and exhausted heat to the cold water reservoir? Number
Answer #1

Solution :-

(A)
Change in entropy between the resrvoirs is

dS = Q*(TH-TC)/(TH*TC)

Q, amount of heat transfered = 36330 J

TH is the temaprature of hot rerservoir = 95 + 273 C = 368 K

TC is the temparature of cool reservoir = 11 + 273 = 284 K


dS = 36330*(368-284)/(368*284)
dS = 29.2j/K

(B)

Work unavailable, = 284 * 29.2J
Work unavailable, = 8292.8 J

(C)

Work done by Carnot Engine is given by -
W/Q = 1 - TC/TH
W = 36330* (1 - 284/368)
W = 8292.72 j